6. The probability of Barcelona winning an upcoming soccer game is 9/10. A represents the event that Barcelona wins the soccer game. The probability of the suns winning an upcoming basketball game is 2/10. B represents the event that the suns win the basketball game. The two events are independent of each other a. (BIA) is the suns winning the upcoming basketball game, given event A (after event A occurs) P(BIA)-
Answer #1

A represent the event that Barcelona wins the soccer game

P(A) = 9/10

B represent the event that the Suns win the basketball game

P(B) = 2/10

A and B are independent events

P(A and B ) = P(A) × P(B)

P(B|A) = P ( A and B ) / P ( A )

P(B|A) = P(A)*P(B) / P(A) = P(B) = 2/10